# Continual Knowledge Adaptation for Reinforcement Learning

### v1

@@ 5.2 Comparison Experiments
Superior model plasticity. As shown in Table 1, our method outperforms SOTA plasticity enhancement techniques dohare2024loss ; abbas2023loss , across all evaluation metrics, including CbpNet and CReLUs. Notably, in SpaceInvaders sequence, our method achieves a significant improvement of 18.30% in average performance (0.8392 $\to$ 0.9928) and an 59.97% increase in forward transfer (0.4844 $\to$ 0.7749) compared to CbpNet. These comprehensive experimental results demonstrate that our approach surpasses existing plasticity enhancement methods, by effectively leveraging historical knowledge for adaptation. This highlights the potential of adaptive processing to improve model plasticity.
@@ 5.4 More Discussions
FT-N wolczyk2021continual 0.3733
ProgNet rusu2016progressive 0.2966
PackNet mallya2018packnet CVPR’2018 0.1900
MaskNet ben2022lifelong TMLR’2023 0.2266
CReLUs abbas2023loss CoLLAs’2023 0.3666
CompoNet malagonself ICML’2024 0.3900
CbpNet dohare2024loss Nature’2024 0.3933
@@
ProgNet rusu2016progressive 0.3125 01938 0.36
PackNet mallya2018packnet CVPR’2018 0.2767 0.1970 0.04
MaskNet ben2022lifelong TMLR’2023 0.0644 -0.0503 0.04
CReLUs abbas2023loss CoLLAs’2023 0.7835 0.7303 0.52
CompoNet malagonself ICML’2024 0.7629 0.7115 0.49
CbpNet dohare2024loss Nature’2024 0.7678 0.7201 0.46
@@
Effectiveness of CKA-RL in Mitigating Catastrophic Forgetting. We use the standard forgetting metric zhang2023replayenhanced , which evaluates performance degradation after training on new tasks. As shown in Table 4, CKA-RL achieves a relatively low forgetting rate (0.45), demonstrating its effective retention of previously learned knowledge while continuing to adapt to new tasks. In contrast, PackNet, which shows the lowest forgetting rate (0.04), sacrifices significant overall performance (0.2767). In contrast, CKA-RL achieves the highest performance (0.7923) and forward transfer (0.7429) across all methods, with a balanced forgetting rate, highlighting its robustness in preventing catastrophic forgetting while maintaining strong performance.
@@ Appendix B More Related Work
Reinforcement Learning (RL) kaelbling1996reinforcement ; 11156142 constitutes a paradigm within machine learning wherein an agent learns to optimize its decision-making process through interaction with an environment. This interaction involves performing actions and receiving consequent feedback, typically in the form of rewards or penalties. The principal learning objective in RL is the maximization of a cumulative reward signal. In contrast to supervised learning, which relies on datasets comprising pre-defined input-output pairs for model training, RL entails an agent acquiring knowledge from the repercussions of its actions, mediated by this reward-penalty mechanism. This iterative, trial-and-error learning process, coupled with its emphasis on sequential decision-making under uncertainty, distinguishes RL from supervised learning methodologies that depend on labeled datasets. Existing reinforcement learning algorithms can be broadly categorized based on whether an explicit model of the environment is learned or utilized, leading to two principal classes: Model-free RL and Model-based RL.
@@
Model-free RL. Model-free RL algorithms enable the agent to learn optimal policies directly from trajectory samples accrued through interaction with the environment, without explicitly constructing an environmental model. Within model-free RL, algorithms are further distinguished by the components they learn, leading to three primary sub-categories: actor-only, critic-only, and actor-critic algorithms. Actor-only algorithms directly learn a policy network, denoted as $\pi_{\theta}(a|s)$, which maps states to actions. This network takes the current state $s_{t}$ as input and outputs the action $a_{t}$. Prominent examples of such algorithms include Reinforce williams1992simple and various policy gradient methods sutton1999policy . Critic-only algorithms, in contrast, focus solely on learning a value function (e.g., state-value or action-value function). Given a state $s_{t}$, the learned value model is used to evaluate all possible actions $a^{\prime}\in A$, and the action $a_{t}$ yielding the maximum estimated value is selected. This category encompasses methods such as Q-learning watkins1989learning . Actor-critic algorithms combine these two approaches by concurrently maintaining and learning both a policy network (the actor) for action selection and a value function model (the critic) for evaluating actions or states. This category includes algorithms such as Deep Deterministic Policy Gradient (DDPG) lillicrap2015continuous , Trust Region Policy Optimization (TRPO) schulman2015trust , Proximal Policy Optimization (PPO) schulman2017proximal , and Asynchronous Advantage Actor-Critic (A3C) mnih2016asynchronous . Notably, PPO has gained considerable traction for training large language models. Recent advancements in this area include GRPO grpo2024 , which employs group‑based advantage estimates within a KL‑regularized loss function to reduce computational overhead and enhance update stability, and DAPO dapo2024 , which utilizes distinct clipping mechanisms and adaptive sampling techniques to improve efficiency and reproducibility during the fine‑tuning of large‑scale models.
@@
Model-based RL. Model-based RL algorithms endeavor to learn an explicit model of the environment, thereby addressing challenges related to sample efficiency. This is because the agent can leverage the learned model for planning and decision-making, reducing the necessity for extensive direct environmental interaction. The learned representation of the environment is commonly termed a ‘world model’. This world model typically predicts the subsequent state $s_{t+1}$ and the immediate reward $r_{t}$ based on the current state $s_{t}$ and the action $a_{t}$ taken. Exemplary model-based RL algorithms include Dyna-Q peng2018deep , Model-Based Policy Optimization (MBPO) janner2019trust , and Adaptation Augmented Model-based Policy Optimization (AMPO) shen2023adaptation .
@@ Appendix C Environments and Tasks
Our continual learning experiments evaluate agents across three complementary task domains designed to test different capabilities: robotic manipulation with continuous control (Meta-World) yu2020meta , and two vision-based Atari environments machado2018revisiting ; bellemare2013arcade emphasizing dynamic decision-making (SpaceInvaders) and sparse-reward navigation (Freeway). This combination spans key RL challenges including high-dimensional state spaces, delayed rewards, procedural variations, and partial observability.
@@ C.1 Meta-World
The robotic manipulation sequence utilizes the Meta-World benchmark yu2020meta , featuring 39-dimensional state observations (encoding arm/object positions) and 4-dimensional continuous actions (arm displacement and gripper torque in $\left[-1,1\right]$). To evaluate continual learning capabilities, we adopt a 20-task sequence consisting of 10 distinct manipulation tasks repeated twice, following the Continual World (CW20) benchmark wolczyk2021continual . This selection covers diverse manipulation skills while maintaining consistent evaluation protocols with prior work. The following lines describe the selected tasks:
@@ Appendix D Implementation Details
To ensure fairness and reproducibility, we build upon the official implementations released by CompoNet malagonself , which provide well-tested baselines for standard reinforcement learning algorithms. Our modifications to the original SAC and PPO codebases are kept minimal: we only substitute the agent definition with our proposed architecture and introduce the corresponding learning mechanisms.
@@
Methods. We compare the CKA-RL with nine SOTA methods. 1) Baseline involves training a randomly initialized neural network for each task, providing a fundamental and essential reference point for comparison. 2) FT-1 (Fine-Tuning Single Model) wolczyk2021continual continuously fine-tunes a single neural network model across all relevant tasks. 3) FT-N (Fine-Tuning with Model Preservation) wolczyk2021continual follows a similar fine-tuning approach but maintains separate model instances for each task to mitigate forgetting. 4) ProgNet rusu2016progressive instantiates a new neural network whenever the task changes, freezing the parameters of the previous modules and adding lateral connections between their hidden layers. 5) PackNet mallya2018packnet stores the parameters to solve every task of the sequence in the same network by building masks to avoid overwriting the ones used to solve previous tasks. 6) MaskNet ben2022lifelong leverages previous score parameters across tasks to learn task-specific score parameters, which are then used to generate masks for the neural network, dynamically adapting the model to new tasks. 7) CReLUs abbas2023loss enhances the agent capability in CRL by concatenating ReLUs, which reduces the incidence of zero activations and addresses the plasticity loss issue in neural networks. 8) CompoNet malagonself introduces a scalable neural network architecture that dynamically composes action outputs from previously learned policy modules, rather than relying on shared hidden layer representations across tasks. 9) CbpNet dohare2024loss uses a variation of back-propagation, continually and randomly reinitializing a small fraction of underutilized units to maintain the plasticity of neural networks.
@@ Appendix E Efficiency of CKA-RL
ProgNet rusu2016progressive 0.4157 -0.0379 0.3757 -0.0075 0.3125 0.1938 0.3680 0.0495
CompoNet malagonself ICML’2024 0.4131 -0.0055 0.9828 0.6963 0.7629 0.7115 0.7196 0.4674

### v2

@@ 5.2 Comparison Experiments
Superior model plasticity. As shown in Table 1, our method outperforms SOTA plasticity enhancement techniques Dohare et al. (2024); Abbas et al. (2023), across all evaluation metrics, including CbpNet and CReLUs. Notably, in SpaceInvaders sequence, our method achieves a significant improvement of 18.30% in average performance (0.8392 $\to$ 0.9928) and an 59.97% increase in forward transfer (0.4844 $\to$ 0.7749) compared to CbpNet. These comprehensive experimental results demonstrate that our approach surpasses existing plasticity enhancement methods, by effectively leveraging historical knowledge for adaptation. This highlights the potential of adaptive processing to improve model plasticity.
@@ 5.4 More Discussions
FT-N Wołczyk et al. (2021) 0.3733
ProgNet Rusu et al. (2016) 0.2966
PackNet Mallya and Lazebnik (2018) CVPR’2018 0.1900
MaskNet Ben-Iwhiwhu et al. (2023) TMLR’2023 0.2266
CReLUs Abbas et al. (2023) CoLLAs’2023 0.3666
CompoNet Malagon et al. (2024) ICML’2024 0.3900
CbpNet Dohare et al. (2024) Nature’2024 0.3933
@@
ProgNet Rusu et al. (2016) 0.3125 01938 0.36
PackNet Mallya and Lazebnik (2018) CVPR’2018 0.2767 0.1970 0.04
MaskNet Ben-Iwhiwhu et al. (2023) TMLR’2023 0.0644 -0.0503 0.04
CReLUs Abbas et al. (2023) CoLLAs’2023 0.7835 0.7303 0.52
CompoNet Malagon et al. (2024) ICML’2024 0.7629 0.7115 0.49
CbpNet Dohare et al. (2024) Nature’2024 0.7678 0.7201 0.46
@@
Effectiveness of CKA-RL in Mitigating Catastrophic Forgetting. We use the standard forgetting metric Zhang et al. (2023b), which evaluates performance degradation after training on new tasks. As shown in Table 4, CKA-RL achieves a relatively low forgetting rate (0.45), demonstrating its effective retention of previously learned knowledge while continuing to adapt to new tasks. In contrast, PackNet, which shows the lowest forgetting rate (0.04), sacrifices significant overall performance (0.2767). In contrast, CKA-RL achieves the highest performance (0.7923) and forward transfer (0.7429) across all methods, with a balanced forgetting rate, highlighting its robustness in preventing catastrophic forgetting while maintaining strong performance.
@@ Appendix B More Related Work
Reinforcement Learning (RL) Kaelbling et al. (1996); Lin et al. (2025) constitutes a paradigm within machine learning wherein an agent learns to optimize its decision-making process through interaction with an environment. This interaction involves performing actions and receiving consequent feedback, typically in the form of rewards or penalties. The principal learning objective in RL is the maximization of a cumulative reward signal. In contrast to supervised learning, which relies on datasets comprising pre-defined input-output pairs for model training, RL entails an agent acquiring knowledge from the repercussions of its actions, mediated by this reward-penalty mechanism. This iterative, trial-and-error learning process, coupled with its emphasis on sequential decision-making under uncertainty, distinguishes RL from supervised learning methodologies that depend on labeled datasets. Existing reinforcement learning algorithms can be broadly categorized based on whether an explicit model of the environment is learned or utilized, leading to two principal classes: Model-free RL and Model-based RL.
@@
Model-free RL. Model-free RL algorithms enable the agent to learn optimal policies directly from trajectory samples accrued through interaction with the environment, without explicitly constructing an environmental model. Within model-free RL, algorithms are further distinguished by the components they learn, leading to three primary sub-categories: actor-only, critic-only, and actor-critic algorithms. Actor-only algorithms directly learn a policy network, denoted as $\pi_{\theta}(a|s)$, which maps states to actions. This network takes the current state $s_{t}$ as input and outputs the action $a_{t}$. Prominent examples of such algorithms include Reinforce Williams (1992) and various policy gradient methods Sutton et al. (1999). Critic-only algorithms, in contrast, focus solely on learning a value function (e.g., state-value or action-value function). Given a state $s_{t}$, the learned value model is used to evaluate all possible actions $a^{\prime}\in A$, and the action $a_{t}$ yielding the maximum estimated value is selected. This category encompasses methods such as Q-learning Watkins (1989). Actor-critic algorithms combine these two approaches by concurrently maintaining and learning both a policy network (the actor) for action selection and a value function model (the critic) for evaluating actions or states. This category includes algorithms such as Deep Deterministic Policy Gradient (DDPG) Lillicrap et al. (2015), Trust Region Policy Optimization (TRPO) Schulman et al. (2015), Proximal Policy Optimization (PPO) Schulman et al. (2017), and Asynchronous Advantage Actor-Critic (A3C) Mnih et al. (2016). Notably, PPO has gained considerable traction for training large language models. Recent advancements in this area include GRPO Shao et al. (2024), which employs group‑based advantage estimates within a KL‑regularized loss function to reduce computational overhead and enhance update stability, and DAPO Yu et al. (2025), which utilizes distinct clipping mechanisms and adaptive sampling techniques to improve efficiency and reproducibility during the fine‑tuning of large‑scale models.
@@
Model-based RL. Model-based RL algorithms endeavor to learn an explicit model of the environment, thereby addressing challenges related to sample efficiency. This is because the agent can leverage the learned model for planning and decision-making, reducing the necessity for extensive direct environmental interaction. The learned representation of the environment is commonly termed a ‘world model’. This world model typically predicts the subsequent state $s_{t+1}$ and the immediate reward $r_{t}$ based on the current state $s_{t}$ and the action $a_{t}$ taken. Exemplary model-based RL algorithms include Dyna-Q Peng et al. (2018), Model-Based Policy Optimization (MBPO) Janner et al. (2019), and Adaptation Augmented Model-based Policy Optimization (AMPO) Shen et al. (2023).
@@ Appendix C Environments and Tasks
Our continual learning experiments evaluate agents across three complementary task domains designed to test different capabilities: robotic manipulation with continuous control (Meta-World) Yu et al. (2020), and two vision-based Atari environments Machado et al. (2018); Bellemare et al. (2013) emphasizing dynamic decision-making (SpaceInvaders) and sparse-reward navigation (Freeway). This combination spans key RL challenges including high-dimensional state spaces, delayed rewards, procedural variations, and partial observability.
@@ C.1 Meta-World
The robotic manipulation sequence utilizes the Meta-World benchmark Yu et al. (2020), featuring 39-dimensional state observations (encoding arm/object positions) and 4-dimensional continuous actions (arm displacement and gripper torque in $\left[-1,1\right]$). To evaluate continual learning capabilities, we adopt a 20-task sequence consisting of 10 distinct manipulation tasks repeated twice, following the Continual World (CW20) benchmark Wołczyk et al. (2021). This selection covers diverse manipulation skills while maintaining consistent evaluation protocols with prior work. The following lines describe the selected tasks:
@@ Appendix D Implementation Details
To ensure fairness and reproducibility, we build upon the official implementations released by CompoNet Malagon et al. (2024), which provide well-tested baselines for standard reinforcement learning algorithms. Our modifications to the original SAC and PPO codebases are kept minimal: we only substitute the agent definition with our proposed architecture and introduce the corresponding learning mechanisms.
@@
Methods. We compare the CKA-RL with nine SOTA methods. 1) Baseline involves training a randomly initialized neural network for each task, providing a fundamental and essential reference point for comparison. 2) FT-1 (Fine-Tuning Single Model) Wołczyk et al. (2021) continuously fine-tunes a single neural network model across all relevant tasks. 3) FT-N (Fine-Tuning with Model Preservation) Wołczyk et al. (2021) follows a similar fine-tuning approach but maintains separate model instances for each task to mitigate forgetting. 4) ProgNet Rusu et al. (2016) instantiates a new neural network whenever the task changes, freezing the parameters of the previous modules and adding lateral connections between their hidden layers. 5) PackNet Mallya and Lazebnik (2018) stores the parameters to solve every task of the sequence in the same network by building masks to avoid overwriting the ones used to solve previous tasks. 6) MaskNet Ben-Iwhiwhu et al. (2023) leverages previous score parameters across tasks to learn task-specific score parameters, which are then used to generate masks for the neural network, dynamically adapting the model to new tasks. 7) CReLUs Abbas et al. (2023) enhances the agent capability in CRL by concatenating ReLUs, which reduces the incidence of zero activations and addresses the plasticity loss issue in neural networks. 8) CompoNet Malagon et al. (2024) introduces a scalable neural network architecture that dynamically composes action outputs from previously learned policy modules, rather than relying on shared hidden layer representations across tasks. 9) CbpNet Dohare et al. (2024) uses a variation of back-propagation, continually and randomly reinitializing a small fraction of underutilized units to maintain the plasticity of neural networks.
@@ Appendix E Efficiency of CKA-RL
ProgNet Rusu et al. (2016) 0.4157 -0.0379 0.3757 -0.0075 0.3125 0.1938 0.3680 0.0495
CompoNet Malagon et al. (2024) ICML’2024 0.4131 -0.0055 0.9828 0.6963 0.7629 0.7115 0.7196 0.4674